# A simple algorithm for output range analysis for deep neural networks

### v3

@@ 5 Simulated Annealing with Boundary Conditions
Let $E=[l_{1},u_{1}]\times\cdots\times[l_{d},u_{d}]\subset\mathbb{R}^{d}$, where $[l_{j},u_{j}]$ are intervals in $\mathbb{R}$. We denote by $\mathcal{R}:\mathbb{R}^{d}\longrightarrow E$, the different combinations of the reflections on $E$, that is, $\mathcal{R}(y)=\big{(}\mathcal{R}(y_{1}),\dots,\mathcal{R}(y_{d})\big{)}$, where
$\mathcal{R}(y_{j})=\begin{cases}l_{j}+\big{[}(y_{j}-l_{j})\mod(u_{j}-l_{j})\big{]},&\text{if }(y_{j}-l_{j})\mod\big{(}2(u_{j}-l_{j})\big{)}\leq u_{j}-l_{j},\\[10.0pt] u_{j}-\big{[}(y_{j}-l_{j})\mod(u_{j}-l_{j})\big{]},&\text{if }(y_{j}-l_{j})\mod\big{(}2(u_{j}-l_{j})\big{)}>u_{j}-l_{j}.\end{cases}$ (5)
@@
We say that ${Q}$ is a generating distribution with Gaussian density function $\rho:E\times\mathbb{R}^{d}\longrightarrow\mathbb{R}^{+}$ and with reflective boundary conditions on $E$, if ${Q}$ is defined as
@@
where $B\subset\mathfrak{B}$, and $\mathcal{B}(E)$ denotes the Borel $\sigma$-algebra on the state space $E$.
@@
for $x\in E$, $y\in\mathbb{R}^{d}$ and $\sigma>0$
@@
$P_{i}(x,B)=\begin{cases}\int\limits_{B}q_{T_{i}}(x,y)\,Q_{i}(x,y)\,\mathrm{d}y&\text{for }x\notin B,\\ \int\limits_{B}q_{T_{i}}(x,y)\,Q_{i}(x,y)\,\mathrm{d}y+\Big{(}1-\int\limits_{E}q_{T_{i}}(x,y)\,Q_{i}(x,y)\,\mathrm{d}y\Big{)}&\text{for }x\in B,\end{cases}$ (9)
@@
This definition aligns with the framework established in Definition 2.4 of Haario and Saksman Haario and Saksman, (1991), which introduces the simulated annealing process in a general state space. Their work provides a rigorous foundation for the transition kernels and the Markov process in this context.
@@
By construction, ${Q}$ is symmetric, that is
@@
for $B\subset\mathfrak{B}$. Furthermore, $\mathcal{F}$ is a continuous function, and added to the fact that $E$ is a compact set, then $M_{\mathcal{F}}$ is a non-empty set and $\mathcal{F}:E\longrightarrow\mathbb{R}$ is uniform continuous. Given these facts, the results in (10) and (11) are direct applications of Theorem 5.1 and Theorem 6.5 in Haario and Saksman, (1991). ∎
@@
The proof is a slight modification of Corollary 5.4 in Haario and Saksman, (1991). ∎
@@ 6 Algorithm Derivation
From Theorem 8 and its Corollary 9, a simple algorithm can easily be derived to solve the range estimation problem $(\mathcal{F},E)$. The main idea is to generate a Markov process $(X_{i})_{i\in\mathbb{N}}$ taking values on $E$, as described in Definition 7, for which its initial state is selected using an arbitrary measure $\mu_{0}$. The following states of the Markov process are selected according to two stages; In the first stage, a new state generated by $Q$ is proposed, and subsequently in the second stage, it is decided to accept or reject that new state according to the probability of acceptance $q$. If the proposed state is outside the domain $E$ we use reflection $\mathcal{R}$ to exchange it for another equally probable state that is inside the domain $E$. It is important to note that the generation of new proposed states using $Q$ is done through $\rho$. Both stages, proposing and accepting new states, are executed repeatedly $N$ times for each temperature level $T$. Theorem 8 guarantees that if we execute this two-stage local recursive process the Markov process $(X_{i})_{i\in\mathbb{N}}$ will converge towards the minimal state corresponding to the value $\mathcal{F}_{\min}$, this occurs for each set temperature level. Finally, Corollary 9 guarantees that if we execute the stages described above for a sequence of temperatures $\{T_{i}\}_{i\in\mathbb{N}}$ that slowly decreases towards zero, then the Markov process $(X_{i})_{i\in\mathbb{N}}$ will converge to the minimal state corresponding to the global minimum that solves our estimation problem. We summarize the entire procedure described above in pseudo-code format in Algorithm 1. In addition, we provide a Python code that is available in https://github.com/Nicerova7/output-range-analysis-for-deep-neural-networks-with-simulated-annealing.
@@ 7.1 The Ackley Function
$f(x_{1},x_{2})=-20\exp\left(-0.2\sqrt{0.5(x_{1}^{2}+x_{2}^{2})}\right)-\exp\left(0.5(\cos 2\pi x_{1}+\cos 2\pi x_{2})\right)+e+20,$ (14)
for which its global minimum point is $f(0,0)=0$, see Figure 1(a). We use this function to obtain a noisy discretized sample $\mathcal{D}=\{(x_{1i},x_{2i},f_{i})\}_{i=1}^{m}\subset E$, where $E=[-4,4]\times[-4,4]$, see 1(b). Furthermore, suppose we want to estimate the function (14) using only the discretized sample of points $\mathcal{D}$. To this end, we use a ResNet as described in the chapter “Residual Neural Networks,” with a total depth of $L=5$ residual blocks. The input layer maps a 2-dimensional input to $H_{0}=128$ neurons. The network contains 4 intermediate residual blocks, each with $H_{l}=256$ neurons for $l=1,\dots,4$, followed by a fifth residual block that reduces the dimensionality back to $H_{5}=128$ neurons before the output layer. The output layer maps the final $128$ neurons to a single output. All layers in the network use the ReLU activation function. The model was initialized with the following setup: Mean Squared Error (MSE) was employed as the loss function, and the Adam optimizer kingma2017adam was used smooth surface generated by ResNet for training with a learning rate of 0.001. The model was trained for 1000 epochs. The fit generated by these ResNet is quite good, resulting in a mean absolute error of $1.0070$ and a mean squared error of $3.0298$ from $1000$ uniform random points generated within the range $[-5,5]\times[-5,5]$. The surface generated by the neural network maintains all the properties of the original function, including its non-convex nature and the existence of many local minima, see Figure 1(c). As mentioned above, we will use this generated response surface to find the global minimum of the Ackley function.
@@
Figure 7.1: Graphs of the experimental evaluation using the Ackley function.
The fact that the generated surface consists of too many local minima constitutes a suitable scenario where our methodology shows its greatest strength. Let us remember that when working with deep neural networks, this scenario is quite recurrent due to the high non-linearity and flexibility of the relationships established by this type of neural network. After training the ResNet $\mathcal{F}$, our main objective is to find the global minimum on the surface generated with the neural network within the domain $E$, denoted as $(E,\mathcal{F}(E))$. To this end, we apply our Algorithm  1 to this case. In figure 7.2, we present the evolution of the values assumed by the Markov process generated through the algorithm and its tendency. We note the reasonably fast convergence towards the global minimum $\mathcal{F}_{\min}=0$. While it is true that there are fluctuations around the trajectory towards convergence, this is due to the highly fluctuating nature of the generated surface $(E,\mathcal{F}(E))$, but there is also a clear and consistent trend towards the global minimum 00.
@@ 7.2 The Drop-Wave Function
for which its global minimum point is $f(0,0)=-1$, see Figure 7.4. In particular, the function is usually evaluated on the square $E=[-5.12,5.12]\times[-5.12,5.12]$. Moreover, suppose we want to estimate the function (15) using only a discretized sample of points $\mathcal{D}=\{(x_{1i},x_{2i},f_{i})\}_{i=1}^{m}\subset E$. For this experiment, we use a total depth of $L=7$ residual blocks. The input layer maps a 2-dimensional input to $H_{0}=128$ neurons. The network contains $L=7$ residual blocks with the following dimensions: $H_{1}=H_{2}=256$, $H_{3}=H_{4}=H_{5}=512$, $H_{6}=256$, and $H_{7}=128$. Finally, The output layer has $128$ neurons. All layers in the network use the ReLU activation function The model was configured with the following setup: The loss function utilized was Mean Squared Error (MSE). Training was performed using the Adam optimizer kingma2017adam with a learning rate of $0.001$ and was performed over 1000 epochs. The fit generated by these ResNets is strong, with a mean absolute error of $0.0221$ and a mean squared error of $0.0010$, based on 1500 points generated within the range $[-5.12,5.12]\times[-5.12,5.12]$. The surface generated by the neural network maintains all the properties of the original function, including its non-convex nature and the existence of many local minima, see Figure 7.5.
@@
This second function is highly complex, and its experiment demonstrates the power of our method. Despite the presence of numerous local minima, our method successfully approximates to the global minimum. This scenario is crucial for solving highly complex problems, which are common in physics and other fields Bezhko_2021; physic_plasma. We applied our Algorithm  1 and obtained the results shown in Figure 7.6. The evolution of values assumed by the Markov process in this instance reveals the full path, where the low values we seek do not always appear at the end of the iteration, but they are still identified. The global minimum, $\mathcal{F}_{\min}=-1$, was successfully found using our method. The trend line clearly illustrates the fluctuations, making them easier to observe. Similar to the previous experiment with the Ackley function, the high fluctuations in the generated surface explain the highly fluctuating path we observed.
@@ 7.3 Multiple Global Minima Function
[…]
. The function is evaluated on the square $E=[-3,3]\times[-3,3]\times[-3,3]$. Then, we want to estimate the function (16) using only a discretized sample of points $\mathcal{D}=\{(x_{1i},x_{2i},,x_{3i},f_{i})\}_{i=1}^{m}\subset E$.
@@
For this function, we use $L=7$ residual blocks, where $H_{0}=128$ represents the neurons in the input layer, $H_{1}=H_{2}=256$, $H_{l}=512$ for $l=3,4,5$, $H_{6}=256$, and $H_{7}=128$, which represents the neurons in the final residual block. The output layer maps the final $128$ neurons to a single output. All layers in the network use the ReLU activation function, except for the final output layer. The model was configured with the following setup: The loss function utilized was Mean Squared Error (MSE). Training was performed using the Adam optimizer with a learning rate of $0.001$ and was performed over 1000 epochs. The fit generated by these ResNets is acceptable, with a mean absolute error of $1.4161$ and a mean squared error of $4.1011$, based on 1500 points generated within the range $[-3,3]\times[-3,3]\times[-3,3]$. The Figure 7.8 illustrates the sample path of our method, showing early convergence towards to one of the global minimum.
@@ 7.3.1 Differences with the Classical Simulated Annealing Algorithm
The Python codes of all the experimental evaluations are available in https://github.com/Nicerova7/output-range-analysis-for-deep-neural-networks-with-simulated-annealing.

### v4

@@ 5 Simulated Annealing with Boundary Conditions
Let $E=[l_{1},u_{1}]\times\cdots\times[l_{d},u_{d}]\subset\mathbb{R}^{d}$, where $l_{j},u_{j}]$ are intervals in $\mathbb{R}$. We denote by $\mathcal{R}:\mathbb{R}^{d}\longrightarrow E$, the different combinations of the reflections on $E$, that is, $\mathcal{R}(y)=\big(\mathcal{R}(y_{1}),\dots,\mathcal{R}(y_{d})\big)$, where
$\mathcal{R}(y_{j})=\begin{cases}l_{j}+\big[(y_{j}-l_{j})\mod(u_{j}-l_{j})\big],&\text{if }(y_{j}-l_{j})\mod\big(2(u_{j}-l_{j})\big)\leq u_{j}-l_{j},\\[10.0pt] u_{j}-\big[(y_{j}-l_{j})\mod(u_{j}-l_{j})\big],&\text{if }(y_{j}-l_{j})\mod\big(2(u_{j}-l_{j})\big)>u_{j}-l_{j}.\end{cases}$
Remark 5 (Intuitive role of the reflection step).
The range estimation problem is inherently posed on bounded input domains—here, a hypercube EE—and any search scheme must therefore keep all iterates within $E$. We employ a cyclic reflection that returns any out-of-box proposal to $E$ via the function $\mathcal{R}$, so that local exploration near the boundary $\partial E$ resembles that of interior points and boundary-driven bias is avoided. Crucially, this boundary mechanism is designed to preserve the assumptions used in our theoretical analysis and ensure the asymptotic convergence of the simulated-annealing chain to the set of global extrema on $E$, see Theorem 9 and Corollary 11. Compared with the general-state-space framework of Haario and Saksman (1991), which is implicitly compatible with boundary handling but does not instantiate an explicit, algorithmic reflection tailored to box-constrained output-range estimation, our contribution makes this reflection function explicit and operational for the present objective.
@@
We say that Q{Q} is a generating distribution, with Gaussian density function $\rho:E\times\mathbb{R}^{d}\longrightarrow\mathbb{R}^{+}$ and with reflective boundary conditions on $E$, if Q{Q} is defined as
@@
where
@@
for $x\in E$, $B\subset\mathfrak{B}(E)$ and $y\in\mathbb{R}^{d}$. Here $\sigma$ is a positive parameter and $\mathfrak{B}(E)$ denote the Borel $\sigma$ -algebra on the state space $E$.
@@
$P_{i}(x,B)=\begin{cases}\int\limits_{B}q_{T_{i}}(x,y)\,Q_{i}(x,y)\,\mathrm{d}y&\text{for }x\notin B,\\ \int\limits_{B}q_{T_{i}}(x,y)\,Q_{i}(x,y)\,\mathrm{d}y+\Big(1-\int\limits_{E}q_{T_{i}}(x,y)\,Q_{i}(x,y)\,\mathrm{d}y\Big)&\text{for }x\in B,\end{cases}$
@@
This definition aligns with the framework established in Definition 2.4 of Haario and Saksman Haario and Saksman (1991), which introduces the simulated annealing process in a general state space. Their work provides a rigorous foundation for the transition kernels and the Markov process in this context.
@@
Remark 10.
The total variation distance between two probability measures $P$ and $Q$ defined on a measurable space $(\Omega,\mathcal{F})$ is given by
$P-Q||_{TV}=\sup_{A\in\mathcal{F}}|P(A)-Q(A)|.$
This measures the largest absolute difference between the probabilities that the two distributions assign to the same event. In the context of our problem, we apply the total variation norm to measure the difference in the distributions of the output range across iterations.
@@
By construction, Q{Q} is symmetric, that is
@@
for $B\subset\mathfrak{B}$. Furthermore, $\mathcal{F}$ is a continuous function, and added to the fact that $E$ is a compact set, then $M_{\mathcal{F}}$ is a non-empty set and $\mathcal{F}:E\longrightarrow\mathbb{R}$ is uniform continuous. Given these facts, the results in (10) and (11) are direct applications of Theorem 5.1 and Theorem 6.5 in Haario and Saksman (1991). ∎
@@
The proof is a slight modification of Corollary 5.4 in Haario and Saksman (1991). ∎
@@ 6 Algorithm Derivation
From Theorem 9 and its Corollary 11, a simple algorithm can easily be derived to solve the range estimation problem $(\mathcal{F},E)$. The main idea is to generate a Markov process $(X_{i})_{i\in\mathbb{N}}$ taking values on $E$, as described in Definition 8, for which its initial state is selected using an arbitrary measure $\mu_{0}$. The following states of the Markov process are selected according to two stages; In the first stage, a new state generated by $Q$ is proposed, and subsequently in the second stage, it is decided to accept or reject that new state according to the probability of acceptance $q$. If the proposed state is outside the domain $E$ we use reflection $\mathcal{R}$ to exchange it for another equally probable state that is inside the domain $E$. It is important to note that the generation of new proposed states using $Q$ is done through $\rho$. Both stages, proposing and accepting new states, are executed repeatedly $N$ times for each temperature level $T$. Theorem 9 guarantees that if we execute this two-stage local recursive process the Markov process $(X_{i})_{i\in\mathbb{N}}$ will converge towards the minimal state corresponding to the value $\mathcal{F}_{\min}$, this occurs for each set temperature level. Finally, Corollary 11 guarantees that if we execute the stages described above for a sequence of temperatures $\{T_{i}\}_{i\in\mathbb{N}}$ that slowly decreases towards zero, then the Markov process $(X_{i})_{i\in\mathbb{N}}$ will converge to the minimal state corresponding to the global minimum that solves our estimation problem. We summarize the entire procedure described above in pseudo-code format in Algorithm 1.
Convergence and cooling schedule.
The theoretical result of the previous section establishes asymptotic convergence under the stated assumptions, compact box $E$, continuity of $\mathcal{F}$, and reflective update via the function $\mathcal{R}$. We do not derive finite-time rates for two complementary reasons. First, in simulated annealing such rates depend on landscape features of $\mathcal{F}$ that, in our black-box setting, are neither observable nor reliably estimable without imposing strong assumptions or accessing internal model information; forcing those assumptions would substantially broaden the scope without adding direct operational value to output-range estimation. Second, the aim of the paper is operational: to provide an effective procedure with asymptotic guarantees and clear usage guidance. In practice, the community often adopts geometric cooling $T_{k+1}=\delta\,T_{k}$ as a pragmatic compromise that accelerates progress empirically, while classical global guarantees are associated with logarithmic schedules, see, e.g., Hajek (1988). The practical guidelines we follow—choosing $\delta$ and tuning the proposal scale to maintain a moderate acceptance rate—are consistent with the classical SA literature and empirical comparisons van Laarhoven and Aarts (1987); Aarts and Korst (1989); Nourani and Andresen (1998); Lundy and Mees (1986). Implementation choices (target acceptance range, proposal scaling, and stopping criterion) are documented in the experimental evaluations section, in line with this framework.
@@ 7.1 The Ackley Function
$f(x_{1},x_{2})=-20\exp\left(-0.2\sqrt{0.5(x_{1}^{2}+x_{2}^{2})}\right)-\exp\left(0.5(\cos 2\pi x_{1}+\cos 2\pi x_{2})\right)+e+20,$
for which its global minimum point is $f(0,0)=0$, see Figure 0(d). We use this function to obtain a noisy discretized sample $\mathcal{D}=\{(x_{1i},x_{2i},f_{i})\}_{i=1}^{m}\subset E$, where $E=[-4,4]\times[-4,4]$, see 0(d). Furthermore, suppose we want to estimate the function (14) using only the discretized sample of points $\mathcal{D}$. To this end, we use a ResNet as described in the chapter “Residual Neural Networks,” with a total depth of $L=5$ residual blocks. The input layer maps a 2-dimensional input to $H_{0}=128$ neurons. The network contains 4 intermediate residual blocks, each with $H_{l}=256$ neurons for $l=1,\dots,4$, followed by a fifth residual block that reduces the dimensionality back to $H_{5}=128$ neurons before the output layer. The output layer maps the final $128$ neurons to a single output. All layers in the network use the ReLU activation function. The model was initialized with the following setup: Mean Squared Error (MSE) was employed as the loss function, and the Adam optimizer Kingma and Ba (2017) was used smooth surface generated by ResNet for training with a learning rate of 0.001. The model was trained for 1000 epochs. The fit generated by these ResNet is quite good, resulting in a mean absolute error of $1.0070$ and a mean squared error of $3.0298$ from $1000$ uniform random points generated within the range $5,5]\times[-5,5]$. The surface generated by the neural network maintains all the properties of the original function, including its non-convex nature and the existence of many local minima, see Figure 0(d). As mentioned above, we will use this generated response surface to find the global minimum of the Ackley function.
@@
(d) Graphs of the experimental evaluation using the Ackley function.
The fact that the generated surface consists of too many local minima constitutes a suitable scenario where our methodology shows its greatest strength. Let us remember that when working with deep neural networks, this scenario is quite recurrent due to the high non-linearity and flexibility of the relationships established by this type of neural network. After training the ResNet $\mathcal{F}$, our main objective is to find the global minimum on the surface generated with the neural network within the domain $E$, denoted as $(E,\mathcal{F}(E))$. To this end, we apply our Algorithm 1 to this case. In figure 7.1, we present the evolution of the values assumed by the Markov process generated through the algorithm and its tendency. We note the reasonably fast convergence towards the global minimum $\mathcal{F}_{\min}=0$. While it is true that there are fluctuations around the trajectory towards convergence, this is due to the highly fluctuating nature of the generated surface $(E,\mathcal{F}(E))$, but there is also a clear and consistent trend towards the global minimum 0.
@@ 7.2 The Drop-Wave Function
for which its global minimum point is f(0,0)=1f(0,0)=-1, see Figure 7.3. In particular, the function is usually evaluated on the square $E=[-5.12,5.12]\times[-5.12,5.12]$. Moreover, suppose we want to estimate the function (15) using only a discretized sample of points $\mathcal{D}=\{(x_{1i},x_{2i},f_{i})\}_{i=1}^{m}\subset E$. For this experiment, we use a total depth of $L=7$ residual blocks. The input layer maps a 2-dimensional input to $H_{0}=128$ neurons. The network contains $L=7$ residual blocks with the following dimensions: $H_{1}=H_{2}=256$, $H_{3}=H_{4}=H_{5}=512$, $H_{6}=256$, and $H_{7}=128$. Finally, The output layer has $128$ neurons. All layers in the network use the ReLU activation function The model was configured with the following setup: The loss function utilized was Mean Squared Error (MSE). Training was performed using the Adam optimizer Kingma and Ba (2017) with a learning rate of $0.001$ and was performed over 1000 epochs. The fit generated by these ResNets is strong, with a mean absolute error of $0.0221$ and a mean squared error of $0.0010$, based on 1500 points generated within the range $5.12,5.12]\times[-5.12,5.12]$. The surface generated by the neural network maintains all the properties of the original function, including its non-convex nature and the existence of many local minima, see Figure 7.4.
@@
This second function is highly complex, and its experiment demonstrates the power of our method. Despite the presence of numerous local minima, our method successfully approximates to the global minimum. This scenario is crucial for solving highly complex problems, which are common in physics and other fields Bezhko et al. (2021); Vijvers et al. (2008). We applied our Algorithm 1 and obtained the results shown in Figure 7.5. The evolution of values assumed by the Markov process in this instance reveals the full path, where the low values we seek do not always appear at the end of the iteration, but they are still identified. The global minimum, $\mathcal{F}_{\min}=-1$, was successfully found using our method. The trend line clearly illustrates the fluctuations, making them easier to observe. Similar to the previous experiment with the Ackley function, the high fluctuations in the generated surface explain the highly fluctuating path we observed.
@@ 7.3 Multiple Global Minima Function
$(x,y,z)\in\{$ $(-1,-1,-1),\,(-1,-1,1),\,(-1,1,-1),$
$(-1,1,1),\,(1,-1,-1),\,(1,-1,1),\,$
$(1,1,-1),\,(1,1,1)\}$
The function is evaluated on the square $E=[-3,3]\times[-3,3]\times[-3,3]$. Then, we want to estimate the function (16) using only a discretized sample of points $\mathcal{D}=\{(x_{1i},x_{2i},,x_{3i},f_{i})\}_{i=1}^{m}\subset E$.
@@
For this function, we use $L=7$ residual blocks, where $H_{0}=128$ represents the neurons in the input layer, $H_{1}=H_{2}=256$, $H_{l}=512$ for $l=3,4,5$, $H_{6}=256$, and $H_{7}=128$, which represents the neurons in the final residual block. The output layer maps the final $128$ neurons to a single output. All layers in the network use the ReLU activation function, except for the final output layer. The model was configured with the following setup: The loss function utilized was Mean Squared Error (MSE). Training was performed using the Adam optimizer with a learning rate of $0.001$ and was performed over 1000 epochs. The fit generated by these ResNets is acceptable, with a mean absolute error of $1.4161$ and a mean squared error of $4.1011$, based on 1500 points generated within the range $3,3]\times[-3,3]\times[-3,3]$. The Figure 7.7 illustrates the sample path of our method, showing early convergence towards to one of the global minimum.
@@ 7.3.1 Differences with the Classical Simulated Annealing Algorithm
The Python codes of all the experimental evaluations are available in this link.
@@ 8 Conclusions
Data availability
The data and source code used in this study are publicly available at the following repository:
GitHub Repository: Output Range Analysis for Deep Neural Networks with Simulated Annealing
This study uses data and code to analyze the output range of deep neural networks based on simulated annealing processes. To ensure the reproducibility of the experiments and facilitate replication, the data and source code are freely accessible.
Source Data
If the data has been previously published, details of the dataset and where it can be accessed should be provided here.
Underlying Data
The repository contains all data used in this study, including input and output files, Python implementation code, and experimental results. The key contents include:
  • $\diamond$
    README.md: Provides an overview of the project, detailing the approach and methodology for output range analysis of deep neural networks using simulated annealing. It includes instructions for reproducing the results and relevant references.
  • $\diamond$
    neural_network.py: Implements the deep neural network architecture used in the study. It defines the DeeperResidualNN class, the loss function (criterion), the optimizer (optimizer), and the training loop for fitting the model to the dataset.
  • $\diamond$
    simulated_annealing.py: Implements the simulated annealing algorithm used to estimate the output range of the neural network. It defines the simulated_annealing function, which optimizes over non-convex surfaces, including parameters such as initial solution, maximum and minimum temperature, cooling rate, number of iterations, intervals, and sigma.
The data and code are freely available without restrictions for use and reproduction.
Competing interests
The authors declare that there are no financial, personal, or professional competing interests that could have influenced the content of this article.
Grant information
The authors declare that no grants were involved in supporting this work.